A 3.7 kg bird traveling 7.99 m/s 50.3 m off the ground swoops down to 8.13 m off the ground. At this point it is traveling 23.2 m/s. What amount of energy is lost to air resistance?


1
Expert's answer
2021-04-19T07:21:26-0400

Solution.

"m=3.7kg;"

"v_1=7.99m\/s;"

"h_1=50.3m;"

"h_2=8.13m;"

"v_2=23.2m\/s;"

"\Delta" "W-?;"

"\Delta W=W1-W_2;"

"W_1=mgh_1+\dfrac{mv_1^2}{2}=m(gh_1+\dfrac{v_1^2}{2});"

"W_2=mgh_2+\dfrac{mv_2^2}{2}=m(gh_2+\dfrac{v_2^2}{2});"

"W_1=3.7kg\sdot(9.81m/s^2\sdot50.3m+\dfrac{(7.99m/s)^2}{2})=1943.8J;"

"W_2=3.7kg\sdot(9.81m/s^2\sdot8.13m+\dfrac{(23.2m/s)^2}{2})=1290.8J;"

"\Delta W=1943.8J-1290.8J=653J;"

Answer: "\Delta W=653J."